### v1

@@ Abstract
This work applies a reduced basis method to study the continuum physics of a finite quantum system—either few or many-body. Specifically, we develop reduced-order models, or emulators, for the underlying inhomogeneous Schrödinger equation and train the emulators against the equation’s bound-state-like solutions at complex energies. The emulators can quickly and accurately interpolate and extrapolate the matrix elements of the Hamiltonian resolvent operator (Green’s function) in a parameter space that encompasses the complex energy plane as well as other real-valued parameters in the Schrödinger equation. The spectra, discretized and compressed as the result of emulation, and the associated resolvent matrix elements (or amplitudes), have the defining characteristics of non-Hermitian quantum mechanics calculations, featuring complex eigenenergies with negative imaginary parts and resolvent matrix elements with branch cuts moved below the real axis in the complex energy plane. Therefore, we now have a method that extracts continuum physics from bound-state-like calculations and emulates them in the input parameter space. Some significant results have previously been presented in a short report (arXiv:2408.03309). Here, we provide details of the study and an in-depth analysis, including how this method could be coupled with existing approaches to compute and emulate bound and continuum states.
@@ I Introduction
In a recent paper Zhang (2024), we reported on a new application of the reduced basis method (RBM) Duguet et al. (2024); Drischler et al. (2023); Melendez et al. (2022) in the study of continuum physics of finite quantum systems. Technical details of the study are provided in this paper. Readers are advised to read the short report first, which could facilitate reading the current paper.
By continuum physics, we mean the part of the spectrum of a Hamiltonian operator where the system can break up into subsystems (i.e., above the system’s lowest threshold) and the states and observables associated with that sector of the spectrum. To be quantitative, let $H(\bm{\theta})$ be the Hamiltonian operator. An important operator is $H$’s resolvent or Green’s function [^1] We are working with time-independent Schrödinger equations in this work.. Its matrix element between two source states $|{S}(\bm{\theta})\rangle$ and $|\tilde{{S}}(\bm{\theta})\rangle$ is named amplitude $\mathcal{A}$, with
@@
The vector $\bm{\theta}$ collects the parameters of $H$, ${S}$, and $\tilde{{S}}$.
$\mathcal{A}$ represents an array of continuum physics observables, such as response functions and scattering amplitudes, depending on the construction of the source states. See Sec. II.1 for the details. The analytical properties of $\mathcal{A}$ in the complex $E$ plane, such as isolated poles and branch cuts (BCTs), are directly connected to the basic features of $H$’s spectrum Newton (1982). This informs an interesting numeric computational framework, called non-Hermitian quantum mechanics (NHQM) Reinhardt (1982); Moiseyev (2011), as discussed in Sec. II.2. Since the approximations of $\mathcal{A}$ produced in a broad class of calculations (including NHQM) could be viewed as rational approximations Trefethen (2023) in terms of the variable $E$, a recent and relevant development on this subject is mentioned in Sec. II.3.
In principle, we can solve the inhomogeneous Schrödinger equation,
@@
To simplify our notation, we assume the $E$ and $\bm{\theta}$ dependence is implicit unless otherwise stated. These equations and $\mathcal{A}$ are the main targets of this study.
@@
We aim to develop RBM-based reduced-order models (ROMs), or emulators, for the solutions of Eqs. (3) and (4) and $\mathcal{A}$ so that we can extrapolate them in the complex $E$ plane and interpolate them in the space of $\bm{\theta}$. Extrapolation in $E$ is a crucial capability, from which $\mathcal{A}$ at real $E$s can be inferred from the solutions at complex $E$s. Solving these equations directly at real $E$s poses severe numeric challenges, but at complex $E$s, it becomes a more feasible bound-state-like calculation. The emulation in $\bm{\theta}$ enables rapid explorations of the continuum physics calculations in the parameter space, another helpful functionality.
@@
As a model-order-reduction (MOR) tool for a parameterized equation system Hesthaven et al. (2015); Quarteroni et al. (2016); Benner et al. (2017a, b, 2015), the RBM first constructs a subspace spanned by the equation’s full (or high-fidelity) solution at a sample of parameter sets, called snapshots, during the so-called offline training stage. Afterward, the equation system is projected into the subspace to form a ROM, which can be used to emulate the solutions and the associated observables in the parameter space. The dimension of the subspace is typically low and scales mildly with the number of parameters Duguet et al. (2024). Consequently, the computing cost for running emulators at the online emulation stage is dramatically reduced compared to simply repeating high-fidelity calculations, e.g., when exploring the parameter space of the calculations.
@@
The basic principle of the RBM was recently rediscovered in nuclear theory as the eigenvector continuation method Frame et al. (2018), where the focus was solving the eigenvalue problem. The RBM-based emulators have gained much attention and further development, including for nuclear-bound states Sarkar and Lee (2021a, 2022); König et al. (2020); Demol et al. (2020); Ekström and Hagen (2019); Demol et al. (2021); Yoshida and Shimizu (2022); Anderson et al. (2022); Giuliani et al. (2023); Yapa et al. (2023), resonant states Yapa et al. (2023, 2024) and general continuum scattering states Furnstahl et al. (2020); Drischler et al. (2021); Melendez et al. (2021); Zhang and Furnstahl (2022); Bai and Ren (2021); Drischler and Zhang (2022); Melendez et al. (2022); Drischler et al. (2023); Bai (2022); Garcia et al. (2023); Odell et al. (2024).
@@
One difference in our work’s RBM aspect is emulating inhomogeneous linear equations with continuous spectra [^2] Ref. Melendez et al. (2022) surveyed different ROMs, including for the inhomogeneous equations. However, the continuous spectrum aspect was not illuminated.. In contrast, most previous quantum physics-related studies have considered emulating a specific eigenstate of a Hamiltonian operator, either a bound, resonance, or scattering state at a real energy $E$. More significantly, this is the first time to consider the complex $E$ plane as part of the parameter space, in addition to the other model input parameters, such as $\bm{\theta}$ in Eqs. (3) and (4). Our RBM formalism is discussed in detail in Sec. III.
@@
If we only consider the $E$ variable, our complex-$E$ emulation (CEE) is superficially similar to the rational Krylov methods Antoulas (2005); Van Beeumen et al. (2017); Peng et al. (2019) applied in studying finite linear equation systems. However, the CEE generalizes the Krylov methods to studying linear systems with continuous spectra. We also gain insights about a potential connection, mentioned throughout this paper, between the CEE and the (near)-optimal rational approximations Trefethen (2023) of a univariate function with branch points. Such a connection does not exist in the case of a linear system with only a discrete spectrum. When including emulation in other real-valued parameters, our study further extends the rational Krylov methods to the case with higher-dimensional parameter spaces; it also generalizes the univariate rational approximation to a multivariate one. We call it CERPE, an abbreviation for “complex-energy real-parameter emulator.” Further discussions on the related works can be found in Sec. III.5
@@
On the physics front, our CEE is a new NHQM method for computing continuum states and observables. This is demonstrated numerically with two and three-body systems in Secs. IV and V. Some analytic understanding of the NHQM aspect of our CEE and the existing NHQM methods, including integration-contour deformation Glöckle (1983), complex scaling of different variants Reinhardt (1982); Moiseyev (2011); Myo et al. (2014); Lazauskas and Carbonell (2011); Lazauskas (2012); Papadimitriou and Vary (2015); Lazauskas (2015); Lazauskas and Carbonell (2020), and Berggren-basis based methods Michel and Płoszajczak (2021); Li et al. (2021); Berggren (1968); Michel et al. (2002, 2003); Id Betan et al. (2003); Hagen et al. (2006); Rotureau et al. (2006); Fossez et al. (2017); Hu et al. (2020); Michel et al. (2022), are presented in Sec. II.2. The generic strategies behind these methods are elaborated in that section using a simple model.
@@
However, our method differs significantly from existing NHQM approaches. The fundamental distinction is in constructing a finite-dimensional non-Hermitian Hamiltonian $H$ matrix—a step we call “non-Hermitization.” [^3] This can also be viewed as breaking time-reversal symmetry in the basis. This difference and its implications are discussed in Sec. III.3.
@@
The CERPE component of this study is also useful for continuum physics studies. Both Hamiltonian spectra and $\mathcal{A}$ can be interpolated, extrapolated, or emulated in the space of $\bm{\theta}$ in the inhomogeneous Schrödinger equations. The functionality of CERPE follows the same argument of existing emulators: they provide efficient interfaces for the users to access computationally expensive calculations with dramatically reduced computing costs Zhang and Furnstahl (2022). For example, with this emulator technology, model calibration and uncertainty quantification, particularly those based on Bayesian statistics, would become feasible for complex models and expensive calculations.
@@
Another new physics insight is concerned with complex-energy (CE) Schlessinger and Schwartz (1966); Schlessinger (1968a, b); McDonald and Nuttall (1969); Uzu et al. (2003); Deltuva and Fonseca (2012, 2013a, 2013b, 2014) and Lorentz integral transformation (LIT) methods Efros (1985a); Efros et al. (1994, 2007); Orlandini et al. (2014); Sobczyk et al. (2021, 2024); Bonaiti et al. (2024). The inhomogeneous Schrödinger equations are also solved at complex energies in the LIT calculations and effectively in the CE calculations. [^4] In the existing implementation of the CE method Uzu et al. (2003); Deltuva and Fonseca (2012, 2013a, 2013b, 2014), the Lippmann-Schwinger and the Faddeev equations are solved for the on- and off-shell scattering amplitudes. However, the wave functions can be computed with those amplitudes and vice versa Schlessinger and Schwartz (1966). Using their procedures, these methods connect the complex-$E$ results to the real-$E$ ones. Our results suggest that these existing calculations can be viewed from the lens of the general NHQM framework. Perhaps more importantly, the CERPE developed here can be applied directly to emulate these existing calculations. The general procedures for achieving this can be found in Sec. VI.
@@
One counterintuitive understanding of computing continuum physics is worth a brief mention. The results shown in this work suggest that the CEE, which approximates $\mathcal{A}$ using a small non-Hermitian $H$-matrix in Eq. (2), produces better results at real energies than the high-fidelity calculations based on a large Hermitian $H$-matrix. What is puzzling is that the two agree numerically in the instances with complex energies, i.e., the emulator is trained by these high-fidelity calculations. In contrast, in the existing emulator studies, emulators are supposed to reproduce high-fidelity calculations, including during extrapolations. We inevitably conclude that the CEE (and thus CERPE) as an extrapolant for $E$ is biased to the physical continuum physics instead of the high-fidelity results based on a discrete spectrum. The non-Hermitization of $H$ plays a key role here. This is further discussed in Sec. VII.
@@
We emphasize that although numerical results are only presented for simple two- and three-body systems with short-range interactions, both CEE and CERPE should work for general finite systems, as the applicability of the RBM method and the working of non-Hermitization are general, without specific reference to the size and the interaction nature of the system. The RBM method requires smooth dependence of the solution on the input parameters, which have been found to hold up in few and many-body studies. Meanwhile, as explained later, the non-Hermitization depends on the spectrum’s continuous nature and the training points’ setup in the complex $E$ plane.
@@
However, we also need to point out that all the numeric calculations here are performed with high accuracy, with relative errors on the order of $10^{-12}$ in the training calculations. Our understanding of the methods developed here is based on such calculations. In practice, the training calculations, although attainable using bound-state methods, can have more significant errors. How the errors impact the performance of the emulator’s extrapolation and ways to stabilize the extrapolation need to be studied in the future.
@@
A summary of the organization of the rest of the paper is as follows. In Sec. II, a general discussion about $H$’s spectrum, its resolvent operator, and their connections to the continuum observables are provided. Recent developments in rational approximation studies are mentioned in light of their relevance in this work. Section III discusses the RBM framework used in this study. Numerical experiments of the CEE and CERPE in both two and three-body systems are discussed and analyzed carefully in Secs. IV and V. In Sec. VI, we discuss the potential couplings between our methods and other calculation methods. In Sec. VII, a summary is provided. The appendices collect some detailed information needed to reproduce the numerical calculations in this work.
@@ II Continuum physics, NHQM methods and rational approximations
Here, we discuss the NHQM approaches for continuum physics and their connections with rational approximations of univariate functions with branch points. Our emulator-based NHQM approach will be briefly motivated towards the end of Sec. II.2, but the formalism is elaborated in Sec. III.
@@ II.1 Scattering and reaction amplitudes
In this work, we focus on the matrix elements of the $H$-resolvent operator with the spatially localized $|S\rangle$ and $|\tilde{S}\rangle$ sources. Per Eq. (5), the matrix elements at real $E$ above thresholds can be computed relying only on the state $|\Psi\rangle$ or $|\tilde{\Psi}\rangle$ at the finite ranges comparable to the spatial size of the sources, without the need for knowing the correct asymptotic behaviors in the wave functions. These matrix elements directly represent or relate to various continuum physics observables. They are discussed separately in the following bullets.
@@
    If $|\tilde{{S}}\rangle=|{S}\rangle=O|\Psi_{\text{bound}}\rangle$ with $O$ as a transition operator and $|\Psi_{\text{bound}}\rangle$ as a bound state, $-\mathrm{Im}\,\mathcal{A}(E)$ is the $O$-induced response function of the $|\Psi_{\text{bound}}\rangle$ state Fetter and Walecka (1972); Efros et al. (2007), when $\mathrm{Im}\,E=0^{+}$. Note the parameters in the sources (i.e., part of $\bm{\theta}$) could be those associated with current operators, such as two-body current parameters (see e.,g., Ref. Pastore et al. (2008)), or the momentum transfer from the probe Walecka (1995).
@@
    The connections between on-shell $\mathcal{A}$ and scattering and reaction $T$-matrices have been discussed in depth in Chap. 16 of Ref. Goldberger and Watson (1964) and Chap. 5 of Ref. Newton (1982) (also see, e.g., Refs. Efros (1985a); Efros et al. (2007)). Suppose $|\Psi_{\text{in}}\rangle$ and $|\Psi_{\text{out}}\rangle$ are the direct product states of cluster internal full wave functions and relative-motion wave functions, properly (anti)symmetrized if needed. Both satisfy the full homogeneous Schrödinger equation $(E-H)|\Psi_{\text{in},\text{out}}\rangle=0$ with real-valued $E$ and large inter-cluster separations. With $E$ being the same as the $E$ variable in the resolvent, $|S\rangle=(H-E)|\Psi_{\text{in}}\rangle$, and similarly $|\tilde{S}\rangle=(H-E)|\Psi_{\text{out}}\rangle$, $\mathcal{A}$ becomes the non-Born term in the corresponding $T$-matrix. The non-Born term is defined here as the full $T$-matrix with the Born term, $\langle\tilde{S}|\Psi_{\text{in}}\rangle=\langle\Psi_{\text{out}}|S\rangle$, subtracted. Note these $|S\rangle$ and $|\tilde{S}\rangle$ states are spatially localized for systems with finite-range interactions because $|\Psi_{\text{in}}\rangle$ and $|\Psi_{\text{out}}\rangle$ satisfy the full Schrödinger equations when the clusters are separated beyond their interaction ranges. Later, we would treat the $E$ variable in the sources and that in the resolvent operator as separate variables if necessary.
@@
    For describing the exclusive transition amplitude induced by a perturbative probe $O$ (e.g., electroweak currents) or its time reversal process, we can use a formalism that mixes up those discussed in the previous two bullet points. I.e., $|S\rangle=O|\Psi_{\text{bound}}\rangle$ and $|\tilde{S}\rangle=(H-E)|\Psi_{\text{out}}\rangle$ or with ${S}$ and $\tilde{{S}}$ exchanged in the assignments. $\mathcal{A}(E)$ gives the full transition amplitude in this situation.
@@
    $\mathcal{A}$ is also an essential and costly component in microscopically computing (nuclear) optical potentials Rotureau et al. (2017); Burrows et al. (2024). The sources here can be chosen to be spatially localized, as the optical potential is intrinsically a quantity defined in the region in which the systems interact.

### v2

@@ Abstract
This work applies a reduced basis method to study the continuum physics of a finite quantum system—either few or many-body. Specifically, I develop reduced-order models, or emulators, for the underlying inhomogeneous Schrödinger equation and train the emulators against the equation’s bound-state-like solutions at complex energies. The emulators rapidly and accurately interpolate and extrapolate the matrix elements of the Hamiltonian resolvent operator (Green’s function) across a parameter space that includes both complex energy and other real-valued physical inputs in the Schrödinger equation. The spectra, discretized and compressed as the result of emulation, and the associated resolvent matrix elements (or amplitudes), have the defining characteristics of non-Hermitian quantum mechanics calculations, featuring complex eigenenergies with negative imaginary parts and branch cuts moved below the real axis in the complex energy plane. Therefore, one now has a method that extracts continuum physics from bound-state-like calculations and emulates those extractions in the input parameter space. Building on a prior Letter [arXiv:2408.03309], this article provides the full theoretical details, a comprehensive analysis of the method’s performance, and a brief discussion of how it can be coupled with existing continuum approaches to perform emulations in their input parameter spaces.
@@ I Introduction
In a recent article [1], I reported on a new application of the reduced basis method (RBM) [2, 3, 4] in the study of continuum physics of finite quantum systems. Technical details of the study are provided in this article. Readers are advised to read the short report [1] first, which could facilitate reading the current article.
By continuum physics, I mean the part of the spectrum of a Hamiltonian operator where the system can break up into subsystems (i.e., above the system’s lowest threshold) and the states and observables associated with that sector of the spectrum. To be quantitative, let $H(\bm{\theta})$ be the Hamiltonian operator. An important operator is $H$’s resolvent or Green’s function that varies with the total energy $E$. Its matrix element between two source states $|{S}(\bm{\theta})\rangle$ and $|\tilde{{S}}(\bm{\theta})\rangle$ is named amplitude $\mathcal{A}$, with
@@
The vector $\bm{\theta}$ collects the parameters of $H$, S{S} and $\tilde{{S}}$.
[…]
In principle, one can solve the inhomogeneous Schrödinger equation,
@@
To simplify the notation, I assume the $E$ and $\bm{\theta}$ dependence is implicit unless otherwise stated. These equations and $\mathcal{A}$ are the main targets of this study.
Solving these equations directly poses severe numeric challenges, when $H$ is Hermitian and $E$ is above continuum thresholds with $\mathrm{Im}\,E\to 0^{+}$ ($0^{-}$), because the solutions are not spatially localized (or integrable) but instead satisfies the outgoing (incoming) boundary conditions in spatial coordinate. However, when away from that energy region (e.g., at complex $E$s), the solutions are bound-state like [17, 18], which are easier to compute.
@@
Here, I aim to develop RBM-based reduced-order models (ROMs), or emulators  [^3] Broadly speaking, emulators, also known as surrogate models, are the tools used to rapidly interpolate and even extrapolate complex calculations or simulations in their parameter spaces [4, 3, 2]., for the solutions of Eqs. (2) and (3) and $\mathcal{A}$ so that one can extrapolate them in the complex $E$ plane and interpolate them in the space of $\bm{\theta}$. Extrapolation in $E$ is a crucial capability, based on which $\mathcal{A}$ at real $E$s can be inferred from the solutions at complex $E$s—as mentioned, the latter are much easier to compute. The $\bm{\theta}$-emulation enables rapid explorations of the continuum physics calculations in the parameter space, another helpful functionality.
@@
As a model-order-reduction (MOR) tool for a parameterized equation system [19, 20, 21, 22, 23], the RBM first constructs a subspace spanned by the equation’s full (or high-fidelity) solution at a sample of parameter sets, called snapshots, during the so-called offline training stage. Afterward, the equation system is projected into the subspace to form a ROM, which can be used to emulate the solutions and the associated observables in the parameter space. The dimension of the subspace is typically low and scales mildly with the number of parameters [2]. Consequently, the computing cost for running emulators at the online emulation stage is dramatically reduced compared to simply repeating high-fidelity calculations, e.g., when exploring the parameter space of the calculations.
@@
The basic principle of the RBM was recently rediscovered in nuclear theory as the eigenvector continuation method [24], where the focus was solving the eigenvalue problem. The RBM-based emulators have gained much attention and further development, including for nuclear-bound states [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35], resonant states [34, 36] and general continuum scattering states [37, 38, 39, 40, 41, 42, 4, 3, 43, 44, 45].
@@
One difference in my work’s RBM aspect is emulating inhomogeneous linear equations with continuous spectra. In contrast, most previous quantum physics-related studies have considered emulating a specific eigenstate of a Hamiltonian operator, either a bound, resonance, or scattering state at a real energy $E$. Note that Ref. [4] surveyed different ROMs, including for the inhomogeneous equations. However, the continuous spectrum aspect was not illuminated. More significantly, this is the first time to consider the complex $E$ plane as part of the parameter space, in addition to the other model input parameters, such as $\bm{\theta}$ in Eqs. (2) and (3). My RBM formalism is discussed in detail in Sec. III.
@@
If one only considers the $E$ variable, the complex-$E$ emulation (CEE) is superficially similar to the rational Krylov methods [46, 47, 48] applied in studying finite linear equation systems. However, the CEE generalizes the Krylov methods to studying linear systems with continuous spectra. One also gains insights about a potential connection, mentioned throughout this paper, between the CEE and the (near-)optimal rational approximations [16] of a univariate function with branch points. Such a connection does not exist in the case of a linear system with only a discrete spectrum. When including emulation in other real-valued parameters, my study further extends the rational Krylov methods to the case with higher-dimensional parameter spaces; it also generalizes the univariate rational approximation to a multivariate one. I call it CERPE, an abbreviation for “complex-energy real-parameter emulator.” Further discussions on the related works can be found in Sec. III.5.
@@
On the physics front, my CEE is a new NHQM method for computing continuum states and observables. This is demonstrated numerically with two and three-body systems in Secs. IV and V. Some analytic understanding of the NHQM aspect of my CEE and the existing NHQM methods, including integration-contour deformation [49], complex scaling of different variants [11, 12, 13, 50, 51, 52, 53, 54], and Berggren-basis based methods [55, 56, 57], are presented in Sec. II.2. The generic strategies behind these methods are elaborated in that section using a simple model.
@@
However, my method differs significantly from existing NHQM approaches. The fundamental distinction is in constructing a finite-dimensional non-Hermitian Hamiltonian $H$ matrix—a step I call “non-Hermitization.” This difference and its implications are discussed in Sec. III.3.
@@
The CERPE component of this study is also useful for continuum physics studies. Both Hamiltonian spectra and $\mathcal{A}$ can be interpolated, extrapolated, or emulated in the space of $\bm{\theta}$ in the inhomogeneous Schrödinger equations. The functionality of CERPE follows the same argument of existing emulators: they provide efficient interfaces for the users to access computationally expensive calculations with dramatically reduced computing costs [40]. For example, with this emulator technology, model calibration and uncertainty quantification, particularly those based on Bayesian statistics, would become feasible for complex models and expensive calculations.
@@
Another new physics insight is concerned with complex-energy (CE) [17, 58, 59, 60, 61, 62, 63, 64, 65] and Lorentz integral transformation (LIT) methods [18, 66, 67, 68, 69, 70, 71]. The inhomogeneous Schrödinger equations are also solved at complex energies in the LIT calculations and effectively in the CE calculations. [^4] In the existing implementation of the CE method [61, 62, 63, 64, 65], the Lippmann-Schwinger and the Faddeev equations are solved for the on- and off-shell scattering amplitudes. However, the wave functions can be computed with those amplitudes and vice versa [17]. Using their procedures, these methods connect the complex-$E$ results to the real-$E$ ones. my results suggest that these existing calculations can be viewed from the lens of the general NHQM framework. Perhaps more importantly, the CERPE developed here can be applied directly to emulate these existing calculations. The general procedures for achieving this can be found in Sec. VI.
@@
One counterintuitive understanding of computing continuum physics is worth a brief mention. The results shown in this work suggest that the CEE, which approximates $\mathcal{A}$ using a small non-Hermitian $H$-matrix in Eq. (1), produces better results at real energies than the high-fidelity calculations based on a large Hermitian $H$-matrix. What is puzzling is that the two agree numerically in the instances with complex energies, i.e., the emulator is trained by these high-fidelity calculations. In contrast, in the existing emulator studies, emulators are supposed to reproduce high-fidelity calculations, including during extrapolations. One inevitably concludes that the CEE (and thus CERPE) as an extrapolant for $E$ is biased to the physical continuum physics instead of the high-fidelity results based on a discrete spectrum. The non-Hermitization of $H$ plays a key role here. This is further discussed in Sec. VII.
@@
I emphasize that although numerical results are only presented for simple two- and three-body systems with short-range interactions, both CEE and CERPE should work for general finite systems, as the applicability of the RBM method and the working of non-Hermitization are general, without specific reference to the size and the interaction nature of the system. The RBM method requires smooth dependence of the solution on the input parameters, which have been found to hold up in few and many-body studies. Meanwhile, as explained later, the non-Hermitization depends on the spectrum’s continuous nature and the training points’ setup in the complex $E$ plane.
@@
However, I also need to point out that all the numeric calculations here are performed with high accuracy, with relative errors on the order of $10^{-12}$ in the training calculations. The presented understanding of my methods is based on such calculations. In practice, the training calculations, although attainable using bound-state methods, can have more significant errors. How the errors impact the performance of the emulator’s extrapolation, and ways to stabilize the extrapolation need to be studied in the future.
@@
A summary of the organization of the rest of the paper is as follows. In Sec. II, a general discussion about $H$’s spectrum, its resolvent operator, and their connections to the continuum observables are provided. Recent developments in rational approximation studies are mentioned in light of their relevance in this work. Section III discusses the RBM framework used in this study. Numerical experiments of the CEE and CERPE in both two and three-body systems are discussed and analyzed carefully in Secs. IV and V. In Sec. VI, I discuss the potential couplings between the methods presented here and other calculation methods. In Sec. VII, a summary is provided. The appendices collect some detailed information needed to reproduce the numerical calculations in this work. Source codes for generating the results of this work can be accessed via the companion website [72].
@@ II Continuum physics, NHQM methods and rational approximations
Here, I discuss the NHQM approaches for continuum physics and their connections with rational approximations of univariate functions with branch points. My emulator-based NHQM approach will be briefly motivated towards the end of Sec. II.2, but the formalism is elaborated in Sec. III.
@@ II.1 Scattering and reaction amplitudes
In this work, I focus on the matrix elements of the $H$-resolvent operator with the spatially localized $|S\rangle$ and $|\tilde{S}\rangle$ sources. Per Eq. (4), the matrix elements at real $E$ above thresholds can be computed relying only on the state $|\Psi\rangle$ or $|\tilde{\Psi}\rangle$ at the finite ranges comparable to the spatial size of the sources, without the need for knowing the correct asymptotic behaviors in the wave functions. These matrix elements directly represent or relate to various continuum physics observables. They are discussed separately in the following bullets.
@@
    If $|\tilde{{S}}\rangle=|{S}\rangle=O|\Psi_{\text{bound}}\rangle$ with $O$ as a transition operator and $|\Psi_{\text{bound}}\rangle$ as a bound state, $\mathrm{Im}\,\mathcal{A}(E)/\pi$ is the $O$-induced response function of the $|\Psi_{\text{bound}}\rangle$ state [73, 67], when $\mathrm{Im}\,E=0^{+}$. Note the parameters in the sources (i.e., part of $\bm{\theta}$) could be those associated with current operators, such as two-body current parameters (see e.,g., Ref. [74]), or the momentum transfer from the probe [75].
@@
    The connections between on-shell $\mathcal{A}$ and scattering and reaction $T$-matrices have been discussed in depth in Chap. 16 of Ref. [76] and Chap. 5 of Ref. [77] (also see, e.g., Refs. [18, 67]). Suppose $|\Phi_{i}\rangle$ and $|\Phi_{f}\rangle$ are the direct product states of cluster internal full wave functions and relative-motion wave functions corresponding to the initial and final states, properly (anti)symmetrized if needed. Both satisfy the full homogeneous Schrödinger equation $(E-H)|\Phi_{i,f}\rangle=0$ with real-valued $E$ and large inter-cluster separations. With $E$ being the same as the $E$ variable in the resolvent, $|S\rangle=(H-E)|\Phi_{i}\rangle$, and similarly $|\tilde{S}\rangle=(H-E)|\Phi_{f}\rangle$, $\mathcal{A}$ becomes the non-Born term in the corresponding $T$-matrix. The non-Born term is defined here as the full $T$-matrix with the Born term, $\langle\tilde{S}|\Phi_{i}\rangle=\langle\Phi_{f}|S\rangle$, subtracted. Note these $|S\rangle$ and $|\tilde{S}\rangle$ states are spatially localized for systems with only two fragments in the initial and final states, because $|\Phi_{i}\rangle$ and $|\Phi_{f}\rangle$ satisfy the full Schrödinger equations when the clusters are separated beyond their interaction ranges. Later, I would treat the $E$ variable in the sources and that in the resolvent operator as separate variables if necessary. Note that these discussions hold for charged systems as well, if Coulomb wave functions are used for the relative motion between fragments.
@@
    For describing the exclusive transition amplitude induced by a perturbative probe $O$ (e.g., electroweak currents) or its time reversal process, one can use a formalism that mixes up those discussed in the previous two bullet points. I.e., $|S\rangle=O|\Psi_{\text{bound}}\rangle$ and $|\tilde{S}\rangle=(H-E)|\Phi_{f}\rangle$ or with S{S} and $\tilde{{S}}$ exchanged in the assignments. $\mathcal{A}(E)$ gives the full transition amplitude in this situation.
@@
    $\mathcal{A}$ is also an essential and costly component in microscopically computing (nuclear) optical potentials [78, 79]. The sources here can be chosen to be spatially localized, as the optical potential is intrinsically a quantity defined in the region in which the systems interact.